41. Solve. 4(7x + 10) = 3(9x – 5) + 20
A) –35
B) 5
C) –5
D) 35
Thank you for your time!!!

Found 2 solutions by tutorcecilia, rcmcc:
Answer by tutorcecilia(2152)

4(7x + 10) = 3(9x – 5) + 20 [Use the distributive property]
28x+40=27x-15+20
28x-27x+40=-15+20
x+40-40=5
x=-35
.
Check by plugging (x=-35) back into the original equation.

Answer by rcmcc(152)

4(7x + 10) = 3(9x – 5) + 20
substitute the terms into the brackets to get the equation
28X+40=27X-15+20
you simplify the equation
28X+40=27X+5
now you subtract 27 from both sides and you get the equation
x+40=5
subtract 40 from each side to isolate x
x=-35
a) is your answer.